find the first five terms in the expansion of (r-2s)^8, can somebody tell me how to find the coeffiecients. Thanks
Answer by stanbon(48535)

find the first five terms in the expansion of (r-2s)^8, can somebody tell me how to find the coeffiecients.
-------------------------------
The form (a-b)^8 is 8C8a^8 + 8C7a^7(-b) + 8C6a^6(-b)^2 + 8C5a^5(-b)^3...
-------------------------------
Your Problem:
(r-(2s))^8= 8C8r^8 + 8C7r^7(-2s) + 8C6r^6(-2s)^2
+ 8C5r^5(-2s)^3 + 8C4r^4(-2s)^4 etc.
-------------------------
The coefficients are:
1st : 1
2nd : 8*(-2) = -16
3rd: 28*(-2)^2 = 28*4 = 112
4th: 56*(-2)^3 = 56*(-8) = -448
5th: I'll leave this to you
==================
cheers,
Stan H.